5.17 grams of CaCl2 was added to 85.0 cm3 of water in a Styrofoam cup calorimeter. The temperature of the water changed from 21.4°C to 27.5 °C.
How many joules energy was required to cause this change? Assume the heat capacity of the system is due only to the water in the system. The specific heat of water is 4.18 J/(g∙°C).
1
Expert's answer
2019-09-23T08:20:24-0400
"q=cm\Delta T""m=\u03c1V"

ρ(H2O) = 1g/cm3

m(H2O)= 1*85.0 = 85 g


"q=4.18\frac{J}{g\times ^\circ C}\times85 g\times (27.5 ^\circ C-21.4 ^\circ C)=2167.33 J =2170 J"

Answer: 2170 J